OpenStudy (anonymous):

Find the exact value of tan5π/4-tanπ/12 all over 1+tan5π/4tanπ/12

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Trig_Identity#Angle_sum_and_difference_identities look up tangent so the expression is equal to tan (5pi/4 - pi/12) --> 5pi/4 - pi/12 = 14pi/12 = 7pi/6 tan(7pi/6) = tan(pi/6) = sqrt3/3

OpenStudy (anonymous):

Cool, thanks that's the same answer i got